A building owner is planning a renovation to an existing building, at an MARR of 12%. As part of this renovation, they have two options:

1) Retain the current HVAC equipment, which has a current market value of $12,000. Its operating costs next year will be $3750, and these costs will increase by 35% each year. Its salvage value will decrease by 20% each year.

2) Obtain a new, more energy-efficient HVAC system at a cost of $18,000. It will have operating costs of $3300 in the first year, which will increase by 30% each year afterwards. Its salvage value will decrease by 24% each year.

Do the following:

A) Determine whether the current system should be replaced now.

B) Determine the replacement strategy considering a 3-year service life.
Answer #1

Given values:

Particulars Current HVAC equipment (In $) Energy-effecient HVAC equipment (In $)
Market value 12000 18000
Operating cost 3750 3300
Gradient 0.35*3750 = 1312.5 0.3*3300 = 990
Salvage value 12000 - (20%*3*12000) = 4800 18000 - (24%*3*18000) = 5040
Service life 3 years 3 years
MARR 12% 12%

let us estimate the equivalent annual worth of cost (EAC) (Since its a cost dominated problem) for both the alternatives.

Current HVAC equipment

EAC = 12000(A/P,12,3) + [3750 + 1312.5(A/G,12,3)] - 4800(A/F,12,3)

Using DCIF Tables

EAC = 12000(0.4163) + (3750 + 1312.5(0.9246)) - 4800(0.2963)

EAC = $8537

Energy-effecient HVAC equipment

EAC = 18000(A/P,12,3) + [3300 + 900(A/G,12,3)] - 5040(A/F,12,3)

Using DCIF tables

EAC = 18000(0.4163) + [3300 + 900(0.9246)] - 5040(0.2963)

EAC = $10132

Based on the calculations the EAC of Current HVAC equipment is lower than Energy-effecient HVAC equipment and hence to be prefered